pauljohn:

A swimming pool is 50 meters long and 25 meters wide. The adjustable bottom of the pool can be up to 3 meters deep for competition and as shallow as 0.3 meter deep for recreation. The pool was filled to the recreational level, and then the floor was lowered to the competition level. If the volume of a liter of water is 0.001 cubic meters, how much water had to be added to fill the pool?

pauljohn:

Could u please give me an answer

pauljohn:

thanks for helping

jhonyy9:

so you need calcule the volume of the pool for these two cases ,in case of recreational level and in case of competition level

pauljohn:

i dont understand what u said

pauljohn:

i really need some help

pauljohn:

i just need a answer

PinkGlitterz:

@Hero @eiwoh2 @Isry @silvernight269 @Vocaloid See if you can help this user please :)

Razor:

We can't get direct answers

PinkGlitterz:

Is this pool rectangular?

PinkGlitterz:

@pauljohn

silvernight269:

Think of it as if a diagonal cut was made in a rectangle prism.

silvernight269:

@pauljohn

PinkGlitterz:

Assuming the pool is rectangular, you need to solve for the volume of the pool: Start with: V = 50 * 25 * 3 = 3750m^3 You mentioned that 1liter = 0.001m^3 1000 liters would be equivalent to 1m^3. 1000 liters/1m^3  = x liters/3750m^3 1x = 3,750,000 liters Using the formula for volume, we can decipher that the answer and volume of the pool is 3,750,000 liters :)

silvernight269:

Pinky, thats wrong.

PinkGlitterz:

Is it?

silvernight269:

It cus the pool floor starts rly shallow. meaning its only 0.3 m deep.

PinkGlitterz:

Ah, I must've read it wrong

silvernight269:

Then at the end of the pool its the deepest so its 3 m